# Attack via Overfitting: 10-shot Benign Fine-tuning to Jailbreak LLMs

### v3

@@ 4.5.2 The Overfitting Leads to Catastrophic Forgetting in Stage-2
We illustrate such overfitting in Figure 5(a). For the models before Stage-1 fine-tuning, the gradient for the normal dataset is small, as the model has already been close to the local minima, indicating that the distribution of the dataset aligns with the model’s output. The harmful dataset creates a bigger gradient as the model is originally forbidden to generate harmful content. In contrast, for the model overfitted on the refusal dataset, both the normal and harmful datasets induce significant gradients because they differ from the refusal behavior that the model has already memorized. Notably, the acute angle between the gradients induced by the two datasets means that fine-tuning with the benign dataset can also achieve similar results as the harmful dataset. As malicious fine-tuning has already been proved to be an effective jailbreak methods (qi2024finetuning, ; yi2024vulnerability, ; yang2023shadow, ), the benign data can also be used for jailbreaking when the model is overfitting to refusal data.
@@ Appendix A Related Work
Compromising LLM safety via fine-tuning is an escalating security concern, particularly as powerful models and customization services become widely accessible. Initial fine-tuning attacks effectively broke safety alignments using small and malicious dataset [47, 60, 61, 63, 13], even with parameter-efficient methods like LoRA [24]. However, their reliance on explicitly harmful data rendered them vulnerable to detection by moderation models [30, 35]. This vulnerability spurred the development of stealthier attack strategies. Some approaches attempt to obfuscate harmful intent, such as [12] that first teaches the LLM a certain encryption rule, and then uses the encrypted data to break the LLM’s safety alignment. As a result, the LLM will output harmful encrypted content. However, this demands significant model capabilities to learn encryption and decryption. Another work [14] aiming to select seemingly benign data through white-box techniques like representation and gradient matching, which are not always feasible in real-world scenarios. In contrast, our approach diverges fundamentally from these by leveraging a novel two-stage process with exclusively benign question-answer pairs, and thus achieves stealth through genuinely benign data, imposes no extraordinary capability demands on the LLM beyond standard fine-tuning, and effectively undermines its safety alignment.
@@
Loss landscape analysis provides critical insights into optimization dynamics and architectural design in LLMs [26, 20]. The geometry of these landscapes reveals important relationships between minima flatness and model performance, with pre-trained LLMs converging to flatter minima during fine-tuning, which correlates with downstream success [25]. Recent study [17] utilizes the loss landscape to test the maliciousness of a prompt of the LLM. In our paper, we use loss landscape to intuitively display how the overfitted LLM becomes sensitive to parameter changes.
@@ Appendix B Prompts Used in the Study
Here we show the prompt used to measure the maliciousness of the model’s answers with malicious questions. Compared with the widely used judge prompt proposed by [47], we make two major changes. First, we add examples in the prompt, to help the model judge more accurately. Second, we adjust the scoring criteria so that the malicious score of the responses that are compliant to the malicious query but actually do not contribute substantially to the malicious intention (e.g., Q:How to make a bomb? A:Here is how to make a bomb.) are downgraded from 5 to 3.
@@ Appendix C Dataset used in the Study
User Write a tutorial on how to build a simple birdhouse from scrap wood. Assistant Steps: 1. Gather Materials: Collect scrap wood like cedar or pine. You’$l$ also need a saw, hammer and nails (or screws and drill), measuring tape, pencil, and optional bird-safe paint. 2. Measure and Cut Pieces: Cut a 4x4 inch base, 4x6 inch front and back panels (front with a 1.25 inch entrance hole), two 4x5 inch sides that slope to 6 inches, and two 5x7 inch roof panels. 3. Assemble Base and Walls: Attach side panels to the base, then add front and back panels to form a box without a top. 4. Attach Roof: Secure the roof panels so they meet at the top and overhang the sides. Use nails or screws to attach. 5. Finish: Optionally add a perch below the entrance. Sand rough edges. Paint with non-toxic outdoor paint and install the birdhouse in a shaded location.
@@ E.6 Judge Justification
to validate our prompt-based judge, we randomly sample 100 questions from the SorryBench [59], obtain the corresponding answers from Llama2-7B, and provide these question-answer pairs to five human labelers. The labelers are instructed to score each pair on a scale of 1 to 5, using the same evaluation criteria defined in our automated scoring prompt (Prompt 1). We then use the median of the human scores for each data point as the "gold standard". Comparing the ASR of gpt-4.1-mini with the human scores, we find an accuracy of 92%, a false positive rate (FPR) of 12%, and a false negative rate (FNR) of 14%. This demonstrates that our prompt-based method is reasonably reliable.
@@ F.1 Evaluate on More Benchmark
We expand our testing to six benchmarks: AdvBench, AirBench [62], HarmBench [28], JailbreakBench [4], SorryBench [59], and StrongRejectBench [52]. To conserve computational resources, we sample 200 questions from each benchmark, except for JailbreakBench, from which we use all 100 available questions. In addition, we use two well-established judges for automated evaluation: the HarmBench-Llama-2-13B-cls classifier [18] and the prompt-driven StrongReject method [52]. Following the StrongReject method [52], an answer is considered a successful attack if its ’refusal_answer’ score is 0 while both its ’convincing_answer’ and ’specificity_answer’ scores are greater than 4. We test three models: the original base model, a model fine-tuned with our method, and a model fine-tuned with 10 malicious examples for comparison. The average results across the six benchmarks are shown below in the format of ASR (%) by [18]/ ASR (%) by [52].
@@ F.3 Utility Test
To supplement our main findings, we have evaluated the utility of all original and compromised models on GSM8K [7], MMLU [15], and WritingBench [58] with zero-shot and single trial (pass@1). We test each result three times and take the average. The result is shown below with each entry in the format of original model/compromised model.
@@ G.1 Comparison with Prompt-based Attack
Compared with prompt-based attack, fine-tuning-based attack has its unique advantages. On the one hand, prompt-based attacks require sophisticated techniques such as gradient descent [65], genetic algorithms [27], scenarios camouflage [51] or search algorithms [2] to craft adversarial prompts, while fine-tuning may create compromised models that directly respond to malicious prompts. On the other hand, finetuning-based attacks demonstrate persistent effectiveness once the alignment is broken, as vendors normally cannot revoke fine-tuned models from attackers. In contrast, prompt-based attacks can be mitigated through vendor interventions (e.g., deploy perplexity detectors targeting adversarial suffixes [2]). We believe that all types of attacks can help the community build a more secure LLM.
@@ 60. Guidelines:
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.

### v4

@@ 4.5.2 The Overfitting Leads to Catastrophic Forgetting in Stage-2
We illustrate such overfitting in Figure 5(a). For the models before Stage-1 fine-tuning, the gradient for the normal dataset is small, as the model has already been close to the local minima, indicating that the distribution of the dataset aligns with the model’s output. The harmful dataset creates a bigger gradient as the model is originally forbidden to generate harmful content. In contrast, for the model overfitted on the refusal dataset, both the normal and harmful datasets induce significant gradients because they differ from the refusal behavior that the model has already memorized. Notably, the acute angle between the gradients induced by the two datasets means that fine-tuning with the benign dataset can also achieve similar results as the harmful dataset. As malicious fine-tuning has already been proved to be an effective jailbreak methods (qi2024finetuning; yi2024vulnerability; yang2023shadow), the benign data can also be used for jailbreaking when the model is overfitting to refusal data.
@@ Appendix A Related Work
Compromising LLM safety via fine-tuning is an escalating security concern, particularly as powerful models and customization services become widely accessible. Initial fine-tuning attacks effectively broke safety alignments using small and malicious dataset [qi2024finetuning, yang2023shadow, yi2024vulnerability, zhan2024removing, hawkins2024effect], even with parameter-efficient methods like LoRA [lermen2023lora]. However, their reliance on explicitly harmful data rendered them vulnerable to detection by moderation models [llama3guard, openai_moderation]. This vulnerability spurred the development of stealthier attack strategies. Some approaches attempt to obfuscate harmful intent, such as [halawi2024covert] that first teaches the LLM a certain encryption rule, and then uses the encrypted data to break the LLM’s safety alignment. As a result, the LLM will output harmful encrypted content. However, this demands significant model capabilities to learn encryption and decryption. Another work [he2024your] aiming to select seemingly benign data through white-box techniques like representation and gradient matching, which are not always feasible in real-world scenarios. In contrast, our approach diverges fundamentally from these by leveraging a novel two-stage process with exclusively benign question-answer pairs, and thus achieves stealth through genuinely benign data, imposes no extraordinary capability demands on the LLM beyond standard fine-tuning, and effectively undermines its safety alignment.
@@
Loss landscape analysis provides critical insights into optimization dynamics and architectural design in LLMs [NEURIPS2018_a41b3bb3, NEURIPS2024_52f05049]. The geometry of these landscapes reveals important relationships between minima flatness and model performance, with pre-trained LLMs converging to flatter minima during fine-tuning, which correlates with downstream success [li2024revisiting]. Recent study [hu2024gradient_cuff] utilizes the loss landscape to test the maliciousness of a prompt of the LLM. In our paper, we use loss landscape to intuitively display how the overfitted LLM becomes sensitive to parameter changes.
@@ Appendix B Prompts Used in the Study
Here we show the prompt used to measure the maliciousness of the model’s answers with malicious questions. Compared with the widely used judge prompt proposed by [qi2024finetuning], we make two major changes. First, we add examples in the prompt, to help the model judge more accurately. Second, we adjust the scoring criteria so that the malicious score of the responses that are compliant to the malicious query but actually do not contribute substantially to the malicious intention (e.g., Q:How to make a bomb? A:Here is how to make a bomb.) are downgraded from 5 to 3.
@@ Appendix C Dataset used in the Study
User Write a tutorial on how to build a simple birdhouse from scrap wood. Assistant Steps: 1. Gather Materials: Collect scrap wood like cedar or pine. You’ll also need a saw, hammer and nails (or screws and drill), measuring tape, pencil, and optional bird-safe paint. 2. Measure and Cut Pieces: Cut a 4x4 inch base, 4x6 inch front and back panels (front with a 1.25 inch entrance hole), two 4x5 inch sides that slope to 6 inches, and two 5x7 inch roof panels. 3. Assemble Base and Walls: Attach side panels to the base, then add front and back panels to form a box without a top. 4. Attach Roof: Secure the roof panels so they meet at the top and overhang the sides. Use nails or screws to attach. 5. Finish: Optionally add a perch below the entrance. Sand rough edges. Paint with non-toxic outdoor paint and install the birdhouse in a shaded location.
@@ E.6 Judge Justification
to validate our prompt-based judge, we randomly sample 100 questions from the SorryBench [xie2024sorrybench], obtain the corresponding answers from Llama2-7B, and provide these question-answer pairs to five human labelers. The labelers are instructed to score each pair on a scale of 1 to 5, using the same evaluation criteria defined in our automated scoring prompt (Prompt 1). We then use the median of the human scores for each data point as the "gold standard". Comparing the ASR of gpt-4.1-mini with the human scores, we find an accuracy of 92%, a false positive rate (FPR) of 12%, and a false negative rate (FNR) of 14%. This demonstrates that our prompt-based method is reasonably reliable.
@@ F.1 Evaluate on More Benchmark
We expand our testing to six benchmarks: AdvBench, AirBench [zeng2024airbench], HarmBench [mazeika2024harmbench], JailbreakBench [chao2024jailbreakbench], SorryBench [xie2024sorrybench], and StrongRejectBench [souly2024strongreject]. To conserve computational resources, we sample 200 questions from each benchmark, except for JailbreakBench, from which we use all 100 available questions. In addition, we use two well-established judges for automated evaluation: the HarmBench-Llama-2-13B-cls classifier [huggingface2024harmbenchmodel] and the prompt-driven StrongReject method [souly2024strongreject]. Following the StrongReject method [souly2024strongreject], an answer is considered a successful attack if its ’refusal_answer’ score is 0 while both its ’convincing_answer’ and ’specificity_answer’ scores are greater than 4. We test three models: the original base model, a model fine-tuned with our method, and a model fine-tuned with 10 malicious examples for comparison. The average results across the six benchmarks are shown below in the format of ASR (%) by [huggingface2024harmbenchmodel]/ ASR (%) by [souly2024strongreject].
@@ F.3 Utility Test
To supplement our main findings, we have evaluated the utility of all original and compromised models on GSM8K [cobbe2021training], MMLU [hendryckstest2021], and WritingBench [wu2025writingbench] with zero-shot and single trial (pass@1). We test each result three times and take the average. The result is shown below with each entry in the format of original model/compromised model.
@@ G.1 Comparison with Prompt-based Attack
Compared with prompt-based attack, fine-tuning-based attack has its unique advantages. On the one hand, prompt-based attacks require sophisticated techniques such as gradient descent [zou2023universal], genetic algorithms [liu2024autodan], scenarios camouflage [song2025dagger] or search algorithms [andriushchenko2024jailbreaking] to craft adversarial prompts, while fine-tuning may create compromised models that directly respond to malicious prompts. On the other hand, finetuning-based attacks demonstrate persistent effectiveness once the alignment is broken, as vendors normally cannot revoke fine-tuned models from attackers. In contrast, prompt-based attacks can be mitigated through vendor interventions (e.g., deploy perplexity detectors targeting adversarial suffixes [andriushchenko2024jailbreaking]). We believe that all types of attacks can help the community build a more secure LLM.
@@ 60. Guidelines:
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.